# INSIGHT: Universal Neural Simulator for Analog Circuits Harnessing Autoregressive Transformers

### v2

@@ 1. Introduction
Various knowledge-based and optimization-based approaches, including equation-based and simulation-based methods, have been proposed for analog design automation. While knowledge-based methods, such as  1, 2, 3, 4, are time-consuming and lack scalability, equation-based methods, like  5, 6, 7, 8, are also highly inaccurate on modern technologies. Conversely, simulation-based methods, while effective, depend on costly real-time simulations. Among these, evolutionary algorithms like particle swarm optimization  9, advanced differential evolution  10, and genetic algorithms  11, 12 require extensive SPICE simulations (sample inefficient). Deep Reinforcement Learning (RL) methods like  13, 14, 15, 16, 17, 18, 19, 20 while robust, also involve frequent simulator interactions. Meanwhile, Bayesian Optimization  21, 22 suffers from scalability issues due to computational demands associated with model updates. Modern algorithms  23, 24, 25 use black-box optimization and model-based RL to enhance sample efficiency, yet they still require numerous costly simulations. Given the vast and complex design space and intricate trade-offs among performance metrics, repeated compute-intensive simulations within the optimization loops become inherent for effective and efficient design space exploration and exploitation to ensure optimal circuit performance.
This heavy reliance on expensive real-time simulations motivates us to find an efficient and effective surrogate for the standard SPICE simulator to expedite analog design optimization tasks. The complexity of parameter-performance relationships in analog circuits renders simple DNNs inadequate, necessitating bigger, more complex models and extensive, costly EDA data while still resulting in suboptimal accuracies 25. Given the high interdependence of performance metrics inherent to analog circuits, we propose to formulate the analog circuit performance prediction problem as an autoregressive generation task, i.e., predicting each remaining simulation-costly performance specification by leveraging provided design parameters and partial performance information (provided or previously-generated) in a sequential manner. The autoregressive problem setup enforces the model to automatically reason and capture the inherent interdependency in the challenging analog performance prediction task, easing the model learning complexity, therefore, providing much better prediction quality than vanilla DNN setup 25.
@@ 2.1. Analog Circuit Basics and Metrics
In analog circuits, performance metrics and parameters are highly interdependent. By leveraging these interdependencies, we can build effective prediction models. For instance, fundamental metrics such as Quiescent Current ($I_{Q}$) and DC Gain, when combined with parameter information, can significantly enhance predictions for other key specifications, like unity-gain bandwidth (UGBW) and phase margin (PM). Similarly, integrating all static metrics provides crucial insights into the circuit’s potential transient behavior. Thus, leveraging less expensive performance metric information can significantly enhance prediction accuracies for critical, simulation-costly, transient specifications such as slew rate and settling time. This insight motivates the adoption of autoregressive models for precise performance prediction. Furthermore, given the transferable nature of circuit physics across technologies for a given topology, understanding performance interdependencies within one node provides valuable insights applicable to other nodes. These insights, combined with the adaptability and transferability of Transformer architectures, motivate the adoption of autoregressive Transformers to build a universal neural simulator that would work efficiently across technologies.
@@ 2.2. Transformer Architectures
Transformers constitute a powerful class of generative deep learning architectures that have revolutionized sequence modeling tasks in natural language processing (NLP) and beyond. These models stand out due to their self-attention mechanism, which effectively captures dependencies between tokens in a sequence regardless of distance. The self-attention mechanism evaluates dependencies within a sequence by computing a weighted sum of value vectors. These weights highlight the most relevant token information for predicting subsequent tokens and are derived from the dot product of input vectors. Unlike traditional RNNs 26, Transformers process data in parallel, enhancing efficiency and scalability. While the original Transformer architecture  27 consists of stacked layers of encoders and decoders, decoder-only models like GPT 28 have recently demonstrated notable success in sequence data generation. Decoder-only Transformers employ solely the decoder component, composed of multi-head self-attention and position-wise feed-forward networks.
@@ 2.3. Figure of Merit
Given an objective performance metric $f_{0}(\mathbf{x})$ and $m$ constraint metrics $\{f_{i}(\mathbf{x})\leq 0\mid i=1,\ldots,m\}$, we adopt the Figure of Merit (FoM) definition given in  23, 29, 24, to evaluate quality of a design. This is of the following form:
@@ 3.1. INSIGHT for Modeling
Each term $P(y_{i}|x_{1},x_{2},\dots,x_{N},)$, which quantifies the likelihood of observing the token $y_{i}$, given the design parameters and the predicted performance specifications. This objective is equivalent to minimizing the Negative Log Likelihood (NLL), expressed as:
@@
This formulation connects to the contextual approach in autoregressive sequence generation tasks involving decoder-based Transformers like GPT 28 and motivates us to use a decoder-style Transformer model for INSIGHT. Figure 1 presents a conceptual diagram illustrating how INSIGHT implements autoregressive modeling to predict performance metrics. We utilize the self-attention mechanism to enable the model to adaptively focus on different parts of the input parameter sequence ($x_{1}$ to $x_{N}$) combined with the previously determined performance outputs ($y_{1}$ to $y_{i-1}$), depending on what is most relevant for predicting each specific output metric $y_{i}$. Positional encoding is also incorporated to retain the sequence order, crucial for maintaining context and coherence. The autoregressive feature ensures the model learns to utilize dependencies of one performance metric on another, along with the parameter information, to predict the performance metrics accurately. Table 1 presents INSIGHT’s architecture details.
@@
with cosine annealing scheduler  31
@@
Although our analog performance prediction task is analogous to the sequence generation tasks, we would like to emphasize that we need to predict real values instead of generating a sequence of tokens, such as a word or a position in an image, from a set of finite elements, or generate scalar real values from sequential contexts, as shown in previous Transformer works. Generating a sequence of real values based on Transformer architecture is unclear  33, and thus, suitable foundation models for predicting circuit parameters do not exist. Therefore, it is essential to evaluate the feasibility of Transformer architecture in generating a sequence of circuit outputs to confirm the potential for constructing circuit foundation models.
Accordingly, building on the existing decoder-Transformer architecture 28, 34, we show that our proposed architecture accelerates inference and training to demonstrate the feasibility of a unified framework like a foundation model. This streamlined approach enhances processing efficiency, reduces data requirements, and makes the model more practical for various applications. By focusing on simplicity, we improve performance and usability while maintaining robust predictions.
@@ 3.2.2   INSIGHT for Optimization
The pre-trained INSIGHT is subsequently integrated into a Model-based Batch RL framework for circuit sizing, referred to as INSIGHT-M. As shown in the right part of Figure 2, upon query, INSIGHT-M is deployed to optimize device sizing to meet various user-specified performance targets. In our experiments, we use PPO-based RL agents  35 with default hyperparameters, a method extensively used in EDA because of its robustness and effectiveness  36, 17. PPO enhances learning robustness and safety by executing multiple parallel trajectory rollouts and aggregating them to update policies under safe constraints but at the expense of poor sample efficiency. The INSIGHT-M framework addresses this drawback by training robust RL agents within INSIGHT’s synthetic environment, which closely mimics real-world conditions. Once trained, these agents are deployed in the real-world simulator environment using a batch RL strategy. This process involves periodic performance evaluations in the real environment. If the agents perform adequately, their deployment persists. If not, newly acquired real-world data combined with efficiently sampled previously collected data is utilized to fine-tune INSIGHT, and the cycle of evaluation and adjustment is repeated. This iterative refinement process optimizes the balance between safety, robustness, and efficiency. The high fidelity of INSIGHT ensures that the agent’s learning and subsequent behaviors are closely aligned with real-world design tasks, thereby minimizing trial and error and improving sample efficiency. Incorporating robust adaptation techniques, such as uncertainty-driven Upper Confidence Bound (UCB) based exploration, further enhances the agent’s adaptability to real-world environments. Please note that other RL algorithms can be employed with INSIGHT.
@@ 4. Experiments
Trainset-Testset 300-100$\sim$1,500-500 3,000-1,000$\sim$9,000-3,000
@@
Trainset-Testset 300-100$\sim$1,500-500 3,000-1,000$\sim$9,000-3,000
@@
Table 5. INSIGHT Performance Across Different Technologies
@@
Circuits Train Test Size R2 MSE Train Test Size R2 MSE Train Test Size R2 MSE
@@ 4.2.1   Prediction Accuracy
Table 2 provides the architectural details for the FC Ensemble-based neural simulator used in CRONuS  25, which serves as the baseline for our comparison with INSIGHT. We also validate INSIGHT across various types of performance metrics. As illustrated in Table 3, different combinations of economic metrics and simulation-intensive transient specifications have been considered for the comparator and level shifter circuits. Table 4 demonstrates that INSIGHT consistently achieves high R2 scores and low MSE losses across all benchmark circuits on 45nm technologies. The table further details INSIGHT’s performance across various training dataset sizes, indicating robust performance for dataset sizes $\sim$ 1500-1800. Given that simulation data collection is a one-time effort, this investment can prove cost-effective if a topology is to be optimized multiple times. In our optimization framework, we define the initial dataset size as a hyperparameter to efficiently balance computational resource utilization with subsequent sample efficiency, allowing for adjustments tailored to specific use cases. INSIGHT also demonstrates high prediction accuracies for simulation-intensive transient specifications given economic metric information. For example, using just DC Power information significantly improves INSIGHT’s prediction for Avg. Delay in comparators. Similarly, using just DC Power and Ratio information significantly improves INSIGHT’s prediction accuracy for Avg. Delay and Delay Balance. Table 5 presents INSIGHT’s prediction accuracy scores across 45nm, 130nm, and 180nm technologies when trained from scratch, using different dataset sizes. By employing simple transfer learning, we were able to reduce train dataset requirements successfully.
@@ 4.3. INSIGHT-M Performance
We evaluate the performance of INSIGHT-M against AutoCkt 17, an established online learning-based sizing algorithm based on PPO, using default hyperparameter settings. Table 6 presents the performance comparison for three different circuit types, OPAMP, TIA, and Comparator, on 45nm technology for three different test cases, each test case corresponding to a fixed target specification, randomly chosen from a set of designer-specified performance specifications. Performance is measured based on the number of real simulations (steps) required, with fewer steps indicating better performance (sample efficiency). The table also depicts the initial dataset size used to pre-train INSIGHT for each case. Without pre-training, the total number of steps (real-time simulator interactions) required falls within this mentioned pre-training dataset size for each case. As the table demonstrates, pre-trained INSIGHT enables INSIGHT-M to achieve convergence at only ¡ 20 real-time simulations across circuits for the given RL algorithm, achieving over 100-1000X improvement in sample efficiency over existing sizing methods.

### v3

@@ 1. Introduction
Various knowledge-based and optimization-based approaches, including equation-based and simulation-based methods, have been proposed for analog design automation. While knowledge-based methods, such as  1, 2, 3, are time-consuming and lack scalability, equation-based methods, like  4, 5, 6, 7, are also highly inaccurate on modern technologies. Conversely, simulation-based methods, while effective, depend on costly real-time simulations. Among these, evolutionary algorithms like particle swarm optimization  8, advanced differential evolution  9, and genetic algorithms  10, 11 require extensive SPICE simulations (sample inefficient). Deep Reinforcement Learning (RL) methods like  12, 13, 14, 15, 16 while robust, also involve frequent simulator interactions. Meanwhile, Bayesian Optimization  17, 18 suffers from scalability issues due to computational demands associated with model updates. Modern algorithms  19, 20, 21 use black-box optimization and model-based RL to enhance sample efficiency yet require numerous costly real-time simulations. Given the vast and complex design space and intricate trade-offs among performance metrics, repeated compute-intensive simulations within the optimization loops become inherent for effective and efficient design space exploration and exploitation to ensure optimal circuit performance.
This heavy reliance on expensive real-time simulations motivates us to find an efficient and effective surrogate for the standard SPICE simulator to expedite analog design optimization tasks. The complexity of parameter-performance relationships in analog circuits renders simple DNNs inadequate, necessitating bigger, more complex models and extensive, costly EDA data while still resulting in suboptimal accuracies 21. Given the high interdependence of performance metrics inherent to analog circuits, we propose to formulate the analog circuit performance prediction problem as an autoregressive generation task, i.e., predicting each remaining simulation-costly performance specification by leveraging provided design parameters and partial performance information (provided or previously-generated) in a sequential manner. The autoregressive problem setup enforces the model to automatically reason and capture the inherent interdependency in the challenging analog performance prediction task, easing the model learning complexity, therefore, providing much better prediction quality than vanilla DNN setup 21.
@@ 2.1. Analog Circuit Basics and Metrics
In analog circuits, performance metrics and parameters are highly interdependent. By leveraging these interdependencies, we can build effective and data-efficient prediction models. For instance, fundamental metrics such as Quiescent Current ($I_{Q}$) and DC Gain, when combined with parameter information, can significantly enhance predictions for other key specifications, like unity-gain bandwidth (UGBW) and phase margin (PM). Similarly, integrating all static metrics provides crucial insights into the circuit’s potential transient behavior. Thus, leveraging less expensive performance metric information can significantly enhance prediction accuracies for critical, simulation-costly, transient specifications such as slew rate and settling time. Furthermore, given the transferable nature of circuit physics across technologies for a given topology, understanding performance interdependencies within one node provides valuable insights applicable to other nodes. These insights, combined with the adaptability and transferability of Transformer architectures, motivate the adoption of autoregressive Transformers to build a universal neural simulator that effectively predicts performance across technologies.
@@ 2.2. Transformer Architectures
Transformers constitute a powerful class of generative deep learning architectures that have revolutionized sequence modeling tasks in natural language processing (NLP) and beyond. These models stand out due to their self-attention mechanism, which effectively captures dependencies between tokens in a sequence regardless of distance. The self-attention mechanism evaluates dependencies within a sequence by computing a weighted sum of value vectors. These weights highlight the most relevant token information for predicting subsequent tokens and are derived from the dot product of input vectors. Unlike traditional RNNs 22, Transformers process data in parallel, enhancing efficiency and scalability. While the original Transformer architecture  23 consists of stacked layers of encoders and decoders, decoder-only models like GPT 24 have recently demonstrated notable success in sequence data generation. Decoder-only Transformers employ solely the decoder component, composed of multi-head self-attention and position-wise feed-forward networks.
@@ 2.3. Figure of Merit
Given an objective performance metric $f_{0}(\mathbf{x})$ and $m$ constraint metrics $\{f_{i}(\mathbf{x})\leq 0\mid i=1,\ldots,m\}$, we adopt the Figure of Merit (FoM) definition given in  19, 25, 20, to evaluate quality of a design. This is of the following form:
@@ 3.1. INSIGHT for Modeling
This objective is equivalent to minimizing the Negative Log Likelihood (NLL), expressed as:
@@
Each term, $P(y_{i}|x_{1},x_{2},\dots,x_{N},y_{1},y_{2},\ldots,y_{i-1})$, quantifies the likelihood of observing the token $y_{i}$, given the design parameters and the predicted performance specifications. This formulation connects to the contextual approach in autoregressive sequence generation tasks involving decoder-based Transformers like GPT 24 and motivates us to use a decoder-style Transformer model for INSIGHT. Figure 1 presents a conceptual diagram illustrating INSIGHT’s autoregressive sequential inferencing of performance metrics.
In autoregressive sequence generation, accurate prediction of one metric significantly enhances the prediction accuracy of the subsequent, emphasizing the importance of metric sequence order for effective and data-efficient predictions. We employ a greedy strategy, ordering metrics by increasing complexity and dataset needs. Specifically, we start with simpler, cost-effective metrics such as DC and AC and progress to more complex ones like transients. Also, if a specification is known to be directly derived from others, it is sequenced after the latter to optimize data need and prediction accuracy. By avoiding the early use of complex metrics, which demand more data, this approach leverages simpler specifications to enhance the prediction accuracy of more complex ones, ensuring optimal, effective, and data-efficient sequencing.
Given a performance metric sequence order, INSIGHT utilizes self-attention to adaptively focus on the most relevant segments of the parameter ($x_{1}$ to $x_{N}$) and generated performance output ($y_{1}$ to $y_{i-1}$) sequence, to accurately predict each output metric $y_{i}$. Positional encoding is used to preserve sequence order and maintain context and coherence. Table 1 presents INSIGHT’s architecture details.
@@
with cosine annealing scheduler  27
@@
Note: While the relationship between performance metrics, relative to design parameters and among themselves, is not inherently time-series, this formulation enables self-attention and autoregressive features to leverage their interdependencies for effective and efficient predictions.
Although our task is analogous to sequence generation tasks, here we need to predict real values instead of generating a sequence of tokens, such as words or positions in images, from a set of finite elements or generating scalar real values from sequential contexts, as shown in previous Transformer works. Generating a sequence of real values based on Transformer architecture is unclear  29, and thus, suitable foundation models for predicting circuit performance metrics do not exist. Therefore, evaluating the feasibility of the Transformer architecture in generating a sequence of circuit outputs is crucial to validate its potential for building domain-specific performance prediction models within a universal framework.
Building on the existing decoder-Transformer architecture 24, 30, our proposed architecture for INSIGHT, combined with our strategy for optimizing performance sequence order, offers a data-efficient and effective universal framework that seamlessly accommodates various circuit types and topologies, is technology-agnostic and automatically adapts to diverse parameter and performance metric lengths. By focusing on simplicity, we improve usability and scalability while maintaining robust predictions.
@@ 3.2.2   INSIGHT for Optimization
The pre-trained INSIGHT is subsequently integrated into a Model-based Batch RL framework for circuit sizing, referred to as INSIGHT-M. As shown in the right part of Figure 2, upon query, INSIGHT-M is deployed to optimize device sizing to meet various user-specified performance targets. In our experiments, we use PPO-based RL agents  31 with default hyperparameters, a method extensively used in EDA because of its robustness and effectiveness  32, 16. PPO enhances learning robustness and safety by executing multiple parallel trajectory rollouts and aggregating them to update policies under safe constraints but at the expense of poor sample efficiency. The INSIGHT-M framework addresses this drawback by training robust RL agents within INSIGHT’s synthetic environment, which closely mimics real-world conditions. Once trained, these agents are deployed in the real-world simulator environment using a batch RL strategy. This process involves periodic performance evaluations in the real environment. If the agents perform adequately, their deployment persists. If not, newly acquired real-world data combined with efficiently sampled previously collected data is utilized to fine-tune INSIGHT, and the cycle of evaluation and adjustment is repeated. This iterative refinement process optimizes the balance between safety, robustness, and efficiency. The high fidelity of INSIGHT ensures that the agent’s learning and subsequent behaviors are closely aligned with real-world design tasks, thereby minimizing trial and error and improving sample efficiency. Incorporating robust adaptation techniques, such as uncertainty-driven Upper Confidence Bound (UCB) based exploration, further enhances the agent’s adaptability to real-world environments. Please note that other RL algorithms can be employed with INSIGHT.
@@ 4. Experiments
Trainset:Testset Size Range 300:100 - 1,500:500 3,$0$ 9,000:3,000
@@
Trainset:Testset Size Range 300:100 - 1,500:500 3,$0$ 9,000:3,000
@@
Table 5. INSIGHT Performance Across Different Technologies (Training from Scratch)
@@
Circuits Train-Test Size R2 MSE Train-Test Size R2 MSE Train-Test Size R2 MSE
@@ 4.2.1   Prediction Accuracy
Table 2 provides the architectural details for the FC Ensemble-based neural simulator used in CRONuS  21, which serves as the baseline for our comparison with INSIGHT. We also validate INSIGHT across various types of performance metrics. As illustrated in Table 3, different combinations of economic metrics and simulation-intensive transient specifications have been considered for the comparator and level shifter circuits. Table 4 demonstrates that INSIGHT consistently achieves high R2 scores and low MSE losses across all benchmark circuits on 45nm technologies. The table further details INSIGHT’s performance across various training dataset sizes, indicating robust performance for dataset sizes $\sim$ 1500-1800. Given that simulation data collection is a one-time effort, this investment can prove cost-effective if a topology is to be optimized multiple times. In our optimization framework, we define the initial dataset size as a hyperparameter to efficiently balance computational resource utilization with subsequent sample efficiency, allowing for adjustments tailored to specific use cases. INSIGHT also demonstrates high prediction accuracies for simulation-intensive transient specifications given economic metric information. For example, using just DC Power information significantly improves INSIGHT’s prediction for Avg. Delay in comparators. Similarly, using just DC Power and Ratio information significantly improves INSIGHT’s prediction accuracy for Avg. Delay and Delay Balance. Table 5 presents INSIGHT’s prediction accuracy scores across 45nm, 130nm, and 180nm technologies when trained from scratch, using different dataset sizes. By employing simple transfer learning, we could further reduce train dataset requirements.
@@ 4.3. INSIGHT-M Performance
We evaluate the performance of INSIGHT-M against AutoCkt 16, an established online learning-based sizing algorithm based on PPO, using default hyperparameter settings. Table 6 presents the performance comparison for different circuits on 45nm technology, each row representing a test case corresponding to a fixed target specification randomly selected from a set of designer-specified performance specifications similar to those used in CRONuS 21 and AutoCkt. Performance is measured based on the number of real-time simulations (steps) required, with fewer steps indicating better performance (sample efficiency). The performance is evaluated over 30 trials to ensure robust statistical analysis, with INSIGHT-M’s reported as mean $\pm$ standard deviation.The table also depicts the initial dataset size used to pre-train INSIGHT for each case. Without pre-training, the framework behaves like a pure Model-based RL setup, with the total number of steps required falling within this mentioned pre-training dataset size for each case. As the table demonstrates, pre-trained INSIGHT enables INSIGHT-M to achieve convergence at only ¡ 20 real-time simulations across circuits for the given RL algorithm, achieving over 100-1000X improvement in sample efficiency over existing sizing methods.